A 0.100M solution of a monoprotic weak acid has a pH of 3.00. What is the pKa of this acid?
Answer #1

PH= 3.00

-log[H+] = 3.00

[H+] = 10^-3.00

[H+]= 10^-3.00 M

Concentration of mono protic acid = 0.100M

for weak acdis

[H+] = square root of KaxC

[H+]^2 = KaxC

Ka = [H+]^2/C = (10^-3)^2/0.1 = 10^-6 x10 = 1.0x10^-5

Ka= 1.0x10^-5

-log(Ka) = -log(1.0x10^-5)

PKa = 5.0